# VEAttack: Downstream-agnostic Vision Encoder Attack against Large Vision Language Models

### v1

@@ 5.4 Ablation studys
Figure 7: VEAttack performance under different attack steps and perturbation budgets on COCO and TextVQA datasets (top x-axis: perturbation budget, bottom x-axis: attack steps).
@@
Following the above observation, we evaluate VEAttack with $t=50$ and $\epsilon=8/255$ on full datasets, comparing the performance and efficiency with traditional white-box attacks. The time costs for traditional white-box methods are estimated based on their evaluated subset size relative to the full dataset, and the Flops count the amount of forward computation. VEAttack surpasses APGD in the image caption task and shows competitive performance in the VQA task. Crucially, compared to the comparison in Figure 2(a), VEAttack further improved attack efficiency, reducing time costs by about 8-fold compared to Ensemble. We hope this efficient implementation demonstrates the potential for practical and scalable adversarial attacks on LVLMs, providing valuable insights for future evaluations.
Table 3: Comparison of the effectiveness and efficiency between VEAttack and other white-box attacks, where "Full Test" refers to the evaluation conducted on the entire dataset.
Attack Full Access Full Test Flops COCO Time (h) Flickr30k Time (h) TextVQA Time (h)
clean 99.3G 115.5 1.33 77.5 0.3 37.1 0.42
APGD croce2020reliable 9.93T 13.1 25.5 9.5 5.2 8.1 22.7
Ensemble schlarmann2023adversarial 9.93T 3.1 41.9 1.2 14.3 0.0 37.8
VEAttack 2.59T 5.5 5.3 4.5 1.1 9.4 4.1
@@
Objective COCO ($\epsilon$$\downarrow$ VQAv2 ($\epsilon$$\downarrow$ POPE ($\epsilon$$\downarrow$
@@
Impact of different attack targets. Table 4 reports the results of using different tokens as attack targets on LLaVa1.5-7B, specifically class token embedding $z_{cls}$ and image token features $z_{v}$. We observe that solely perturbing $z_{cls}$ results in relatively minor attack performance across all datasets, indicating its limited ability to effectively disrupt LVLMs. In contrast, simultaneously attacking both $z_{cls}$ and $z_{v}$ leads to more noticeable attack performance, demonstrating the superior effectiveness of $z_{v}$ over $z_{cls}$. Notably, the greatest attack performance is achieved when attacking only $z_{v}$, indicating that $z_{v}$ can sufficiently represent visual information. These results are consistent with our theoretical analysis in Section 4.2, confirming that perturbations on $z_{v}$ propagate more effectively to the downstream, resulting in more effective attacks.
@@
Measurement COCO ($\epsilon$$\downarrow$ VQAv2 ($\epsilon$$\downarrow$ POPE ($\epsilon$$\downarrow$
@@ 6 Conclusion
We propose VEAttack, a novel white-box attack targeting the vision encoder of LVLMs, balancing reliance on task-specific gradients to overcome the task-specificity and high computational costs of traditional attacks. By minimizing cosine similarity between clean and perturbed visual features, VEAttack ensures the generalization across tasks like image captioning and visual question answering, achieving performance degradation of $94.5\%$ and $75.7\%$, respectively. It reduces computational overhead 8-fold compared to ensemble methods and eliminates task and label dependencies via downstream-agnostic perturbations. Our experiments reveal key LVLM vulnerabilities, including hidden layer variations, token attention differentials, Möbius band in transfer attack, and low sensitivity to attack steps. We believe the insights of VEAttack will drive further research into LVLM robustness, fostering advanced defense mechanisms for secure multimodal AI systems.
@@
Limitations. A limitation of this work is that we have not yet investigated defense mechanisms to counter VEAttack. Although more robust vision encoders may improve LVLM defenses, significant performance degradation persists under white-box conditions of the vision encoder. And mitigating transfer attacks from adversarial samples targeting such robust vision encoders is unresolved. Addressing these issues is crucial for ensuring the secure deployment of LVLMs.
Appendix A Proof of Proposition 1
@@
The cross-modal alignment mechanism $f_{A}:\mathbb{R}^{n_{v}\times d_{v}}\to\mathbb{R}^{n_{v}\times d_{m}}$ is a linear projection layer with weight $W_{a}$, where $z_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$ is the output of vision encoder. After our image tokens attack on vision encoder CLIP, the difference of output features during attack will have a lower bound, which can be denoted as $\Delta z_{v}=\tilde{z}_{v}-z_{v}$ and $\|\Delta z_{v}\|_{F}\geq\Delta$, where $\tilde{z}_{v}$ is the output of vision encoder after adversarial attack. So the aligned feature can be formulated as:
@@
Appendix B Proof of Proposition 2
The vision encoder CLIP-ViT processes an input image by dividing it into $n_{v}$ image tokens $z_{v}$ and a class embedding $z_{cls}$. In Layer 0, the initial token can be denoted as $T_{0}=[z_{cls}^{0};z_{v}^{0,1};\dots;z_{v}^{0,n_{v}}]\in\mathbb{R}^{(n_{v}+1)\times d_{v}}$, where $z_{v}^{0,i}$ represents the embedding of the $i$-th patch. The ViT comprises $L$ layers, applying a self-attention mechanism followed by a residual connection:
@@ Appendix B Proof of Proposition 2
[…]
Define $M_{i}=A_{i,0}(A_{0,0}W_{V}+I)^{T}W_{V}+\sum_{j=1}^{n_{v}}A_{i,j}(A_{0,j}W_{V})^{T}W_{V}+(A_{0,i}W_{V})^{T}$, then we can simplify the perturbation as $\Delta z_{v}^{1,i}=-M_{i}\Delta z_{cls}^{1}$. For all image tokens, we can aggregate as:
@@
If the feature of attack objective is $z_{cls}$, following the derivation in Section A, we have:
@@
This matches the second claim of Proposition 2. When extended to $L$-layer CLIP, the above proof occurs between $L$ and $L-1$ layers, where the propagation is in the form of $[z_{cls}^{L}+\Delta z^{L}_{cls};z_{v}^{L}]\rightarrow[z_{cls}^{L-1}+\Delta z^{L-1}_{cls};z_{v}^{L-1}+\Delta z^{L-1}_{v}]\rightarrow[z_{cls}^{L};z_{v}^{L}+\Delta z^{L}_{v}]$. Before $L-1$ layer, the propagation follows the $[z_{cls}^{i}+\Delta z^{i}_{cls};z_{v}^{i}+\Delta z^{i}_{v}]\rightarrow[z_{cls}^{i-1}+\Delta z^{i-1}_{cls};z_{v}^{i-1}+\Delta z^{i-1}_{v}]$ propagation method. Since we have assumed the amplitude equivalence before and after the propagation, we simplify the propagation of the first $L-1$ layer as having the same amplitude. Therefore, our conclusion still holds.
Appendix C More results on multi-task transfer attack
In Figure 2(b) of the main text, we initially demonstrated the transfer attack between COCO Caption and VQAv2 datasets, representing two distinct downstream tasks, and confirmed that VEAttack effectively achieves downstream-agnostic attacks across these tasks. Table 6 provides a more detailed and comprehensive analysis of multi-task transfer attack performance. We use "Anytask" to denote that VEAttack targets any task, such as COCO Caption or VQA, as it solely attacks the vision encoder, enabling downstream-agnostic performance that remains consistent across different tasks. For APGD, we underline the performance of transfer attacks between tasks, revealing that these reductions (e.g., $49.4\%$ for VQAv2 to COCO Caption, and $29.2\%$ for VQAv2 to OK_VQA) are consistently less severe than those achieved by VEAttack (e.g., $93.9\%$ for COCO Caption, and $62.8\%$ for OK_VQA). This highlights the limitations of traditional white-box attacks like APGD in multi-task LVLM scenarios, while further validating the effectiveness and robustness of VEAttack in achieving significant performance degradation across diverse tasks.
Table 6: Transfer Attack Performance of APGD and VEAttack Between Captioning and VQA Tasks. The underlined values in APGD represent transfer attack results and diagonal values indicate direct white-box attack performance.
@@ Appendix C More results on multi-task transfer attack
Appendix D More evaluation metrics on image captioning
Table 7 presents a comprehensive evaluation of the VEAttack method on image captioning across the COCO and Flickr30k datasets using the LLaVa1.5-7B model. Compared to the clean baselines, VEAttack consistently demonstrates the most significant performance degradation on COCO captioning across all evaluation metrics. Specifically, $\text{BLEU}@4$, which focuses on measuring caption precision, shows a substantial drop of $86.7\%$ on the COCO dataset, indicating the highly effective nature of our VEAttack. ROUGE_L, which evaluates recall and sequence overlap, declines moderately $45.8\%$, suggesting that while the attack disrupts overall performance, some structural coherence in the captions remains. SPICE, focusing on semantic propositional content, drops $87.1\%$, indicating a broad degradation in meaningful content representation, reinforcing the effectiveness of VEAttack.
@@ Appendix D More evaluation metrics on image captioning
Dataset Attack $\text{BLUE}@4$ METEOR ROUGE_L CIDEr SPICE
@@
Appendix E Evaluation on image classification datasets

### v2

@@ 5.4 Ablation studys
Figure 8: VEAttack performance under different steps and budgets on COCO and TextVQA datasets (top x-axis: perturbation budget, bottom x-axis: attack steps).
@@
Following the above observation, we evaluate VEAttack with $t=50$ and $\epsilon=8/255$ on full datasets, comparing the performance and efficiency with white-box and gray-box attacks. The time costs for traditional white-box methods are estimated based on their evaluated subset size relative to the full dataset, and the Flops count the computation of forward once. VEAttack surpasses APGD in the image caption task and shows competitive performance in the VQA task. Crucially, compared to the comparison in Fig. 2(a), VEAttack further improved attack efficiency, reducing time costs by about 8-fold compared to white-box Ensemble attack, and about 13-fold compared to gray-box VT-Attack with 1000 iterations. We hope this efficient implementation demonstrates the potential for practical and scalable adversarial attacks on LVLMs, providing valuable insights for future evaluations.
Table 3: Comparison of the effectiveness and efficiency between VEAttack and other white-box and gray-box attacks. Flops count the computation of forward once.
Attack Version Flops COCO Time (h) Flickr30k Time (h) TextVQA Time (h)
clean None 99.3G 115.5 1.33 77.5 0.3 37.1 0.42
APGD Croce and Hein (2020) White-box 9.93T 13.1 25.5 9.5 5.2 8.1 22.7
Ensemble Schlarmann and Hein (2023) White-box 9.93T 3.1 41.9 1.2 14.3 0.0 37.8
VT-Attack Wang et al. (2024c) Gray-box 3.04T 12.2 65.0 12.3 13.3 10.0 64.8
VEAttack ($\epsilon=4/255,t=100$) Gray-box 2.59T 7.1 8.5 6.3 1.7 10.1 6.7
VEAttack ($\epsilon=8/255,t=50$) Gray-box 2.59T 5.5 5.3 4.5 1.1 9.4 4.1
@@
[…]
@@
Impact of different attack targets. Table 4 shows the results of using different tokens as attack targets on LLaVa1.5-7B, specifically class token embedding $z_{cls}$ and patch token features $z_{v}$. We observe that solely perturbing $z_{cls}$ results in relatively minor attack performance across all datasets, indicating its limited ability to effectively disrupt LVLMs. In contrast, simultaneously attacking both $z_{cls}$ and $z_{v}$ leads to more noticeable attack performance, demonstrating the superior effectiveness of $z_{v}$ over $z_{cls}$. Notably, the greatest attack performance is achieved when attacking only $z_{v}$, indicating that $z_{v}$ can sufficiently represent visual information. These results are consistent with our theoretical analysis in Section 4.2, confirming that perturbations on $z_{v}$ propagate more effectively to the downstream, resulting in effective attacks.
@@
[…]
@@ 6 Conclusion
We propose VEAttack, a novel gray-box attack targeting solely the vision encoder of LVLMs, balancing reliance on task-specific gradients to overcome the task-specificity and high computational costs. By minimizing cosine similarity between clean and perturbed visual features under limited perturbation sizes, VEAttack ensures the generalization across tasks like image captioning and visual question answering, achieving performance degradation of $94.5\%$ and $75.7\%$, respectively. It reduces computational overhead 8-fold compared to white-box Ensemble attack and 13-fold compared to gray-box VT-Attack. Our experiments reveal key LVLM vulnerabilities, including hidden layer variations, token attention differentials, Möbius band in transfer attack, and low sensitivity to attack steps. We believe the insights of VEAttack will drive further research into LVLM robustness, fostering advanced defense mechanisms for secure multimodal AI systems.
@@
Limitations. A limitation of this work is that we have not yet investigated defense mechanisms to counter VEAttack. Although more robust vision encoders may improve LVLM defenses, significant performance degradation persists under gray-box conditions of the vision encoder. And mitigating transfer attacks from adversarial samples targeting such robust vision encoders is unresolved. Addressing these issues is crucial for ensuring the secure deployment of LVLMs.
Reproducibility statement
We aim to make VEAttack fully reproducible. The algorithmic specification and assumptions are given in Section  4.1 to Section 4.2) (including Eqs. (4)–(7)), with the overall pipeline summarized in Fig. 4, while formal guarantees and complete proofs appear in Appendix B and C. Experimental settings, including model variants, datasets, evaluation metrics, and the gray‑box attack settings, are introduced in Section 5.1, while primary results and ablations are reported in Section 5.2 to Section 5.4 (experimental results are shown in Tables 1 to Table 5 and Figs. 2 to 8). Additional materials that aid verification include multi‑task transfer results (Appendix D, Table 6), effects and costs of joint LLM (Appendix E, Table 7), extended captioning metrics (Appendix F, Table 8), Image-Text Retrieval task (Appendix G, Table 10), transfer attack from larger specific vision encoders (Appendix I, Table 13), classification evaluations (Appendix J, Table 14), and qualitative analyses (Appendix K, Figs. 10 to 13). To facilitate replication, our supplementary materials contain an anonymous code file that implements VEAttack, along with a README and run scripts that regenerate all our results.
Ethics statement
We affirm adherence to the Code of Ethics throughout this work. Our study investigates adversarial robustness of LVLMs in a research‑only setting and involves no human‑subjects experiments, user studies, or collection of personal data. All experiments use publicly available datasets and pretrained models (e.g., COCO, Flickr30k, TextVQA, VQAv2, LLaVA, OpenFlamingo, miniGPT‑4, mPLUG‑Owl2, Qwen‑VL, and CLIP), following their licenses and prescribed evaluation protocols, and all sources are cited in the paper. We recognize the dual‑use risk inherent in security‑oriented research on adversarial attacks. We intend to expose vulnerabilities to improve robustness: we study an untargeted, vision‑encoder–only attack under small perturbation budgets, report results transparently, and release any accompanying code solely to support reproducibility and robustness evaluation, not real‑world misuse.
Appendix A Ethical Statement and Responsible Disclosure
As Large Vision-Language Models (LVLMs) are increasingly deployed in safety-critical applications, investigating their adversarial vulnerabilities is essential for building resilient systems. This work introduces VEAttack, a method designed to expose inherent weaknesses in the vision encoder component of LVLMs. While our research demonstrates effective attack capabilities, our primary motivation is to advance the community’s understanding of multimodal robustness and to facilitate the development of stronger defense mechanisms.
Responsible Release. To mitigate potential misuse, we adhere to responsible disclosure principles. We will release the source code and generated adversarial examples strictly for research purposes under a restrictive license that prohibits malicious use. Furthermore, we have consciously excluded any functionality that targets specific harmful or offensive content generation, focusing solely on benign benchmarks to demonstrate technical vulnerabilities without generating toxic outputs.
Safety Implications and Mitigation. The Möbius Band phenomenon identified in this study highlights a critical safety paradox: robust vision encoders can inadvertently serve as potent sources for transfer attacks. This finding underscores the urgent need for holistic defense strategies that go beyond standard adversarial training. We advocate for future research to focus on: 1) developing transfer-resistant vision encoders that decouple robustness from attack transferability, and 2) implementing multi-stage verification protocols in LVLM pipelines to detect and filter adversarial visual tokens before they propagate to the language model.
By publicly documenting these vulnerabilities and their mechanistic underpinnings, we aim to provide the necessary groundwork for the research community to preemptively secure future multimodal systems against emerging gray-box threats.
Appendix B Proof of Proposition 1
@@
The cross-modal alignment mechanism $f_{A}:\mathbb{R}^{n_{v}\times d_{v}}\to\mathbb{R}^{n_{v}\times d_{m}}$ is a linear projection layer with weight $W_{a}$, where $z_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$ is the output of vision encoder. After our visual tokens attack on vision encoder CLIP, the difference of output features during attack will have a lower bound, which can be denoted as $\Delta z_{v}=\tilde{z}_{v}-z_{v}$ and $\|\Delta z_{v}\|_{F}\geq\Delta$, where $\tilde{z}_{v}$ is the output of vision encoder after adversarial attack. So the aligned feature can be formulated as:
@@
Appendix C Proof of Proposition 2
The vision encoder CLIP-ViT processes an input image by dividing it into $n_{v}$ patch tokens $z_{v}$ and a class embedding $z_{cls}$. In Layer 0, the initial token can be denoted as $T_{0}=[z_{cls}^{0};z_{v}^{0,1};\dots;z_{v}^{0,n_{v}}]\in\mathbb{R}^{(n_{v}+1)\times d_{v}}$, where $z_{v}^{0,i}$ represents the embedding of the $i$-th patch. The ViT comprises $L$ layers, applying a self-attention mechanism followed by a residual connection:
@@ Appendix C Proof of Proposition 2
$\begin{aligned} \Delta z_{v}^{1,i}&=-A_{i,0}\left[(A_{0,0}W_{V}+I)^{T}\Delta z_{cls}^{1}W_{V}\right]-\sum_{j=1}^{n_{v}}A_{i,j}[(A_{0,j}W_{V})^{T}\Delta z_{cls}^{1}W_{V}]-(A_{0,i}W_{V})^{T}\Delta z_{cls}^{1}\\ &=-\left[A_{i,0}(A_{0,0}W_{V}+I)^{T}W_{V}+\sum_{j=1}^{n_{v}}A_{i,j}(A_{0,j}W_{V})^{T}W_{V}+(A_{0,i}W_{V})^{T}\right]\Delta z_{cls}^{1}\end{aligned}$
Define $M_{i}=A_{i,0}(A_{0,0}W_{V}+I)^{T}W_{V}+\sum_{j=1}^{n_{v}}A_{i,j}(A_{0,j}W_{V})^{T}W_{V}+(A_{0,i}W_{V})^{T}$, then we can simplify the perturbation as $\Delta z_{v}^{1,i}=-M_{i}\Delta z_{cls}^{1}$. For all patch tokens, we can aggregate as:
@@
If the feature of attack objective is $z_{cls}$, following the derivation in Section B, we have:
@@
This matches the second claim of Proposition 2. When extended to $L$-layer CLIP, the above proof occurs between $L$ and L1L-1 layers, where the propagation is in the form of $z_{cls}^{L}+\Delta z^{L}_{cls};z_{v}^{L}]\rightarrow[z_{cls}^{L-1}+\Delta z^{L-1}_{cls};z_{v}^{L-1}+\Delta z^{L-1}_{v}]\rightarrow[z_{cls}^{L};z_{v}^{L}+\Delta z^{L}_{v}]$. Before L1L-1 layer, the propagation follows the $z_{cls}^{i}+\Delta z^{i}_{cls};z_{v}^{i}+\Delta z^{i}_{v}]\rightarrow[z_{cls}^{i-1}+\Delta z^{i-1}_{cls};z_{v}^{i-1}+\Delta z^{i-1}_{v}]$ propagation method. Since we have assumed the amplitude equivalence before and after the propagation, we simplify the propagation of the first L1L-1 layer as having the same amplitude. Therefore, our conclusion still holds.
Appendix D More results on multi-task transfer attack
In Fig. 2(b) of the main text, we initially demonstrated the transfer attack between COCO Caption and VQAv2 datasets, representing two distinct downstream tasks, and confirmed that VEAttack effectively achieves downstream-agnostic attacks across these tasks. Table 6 provides a more detailed and comprehensive analysis of multi-task transfer attack performance. We use ”Anytask” to denote that VEAttack targets any task, such as COCO Caption or VQA, as it solely attacks the vision encoder, enabling downstream-agnostic performance that remains consistent across different tasks. For APGD, we underline the performance of transfer attacks between tasks, revealing that these reductions (e.g., $49.4\%$ for VQAv2 to COCO Caption, and $29.2\%$ for VQAv2 to OK_VQA) are consistently less severe than those achieved by VEAttack (e.g., $93.9\%$ for COCO Caption, and $62.8\%$ for OK_VQA). This highlights the limitations of traditional white-box attacks like APGD in multi-task LVLM scenarios, while further validating the effectiveness and robustness of VEAttack in achieving significant performance degradation across diverse tasks.
Table 6: Transfer Attack Performance of APGD and VEAttack Between Captioning and VQA Tasks. The underlined values in APGD represent transfer attack results, and diagonal values indicate direct white-box attack performance.
@@ Appendix D More results on multi-task transfer attack
Appendix E Effects and cost of VEAttack with LLM
To further demonstrate whether the effect of attacking the vision encoder is sufficient, we add a joint attack on LLM to the current VEAttack, where we supervised all the hidden states of LLM outputs. The results in the Table 7 demonstrate a further enhancement in attack effectiveness on the Flickr30k dataset when attacking jointly the vision encoder and LLM, with a performance gain of 0.6. However, this improvement comes at a significant computational cost, increasing runtime by $278\%$. Through this observation, LLM shows a limited effect in the attack but highly increases the computational cost, further validating the necessity of focusing on vision encoder attacks as proposed in our framework.
Table 7: Comparison of VEAttack and VEAttack+LLM backbone attack on LLaVa1.5-7B.
Method Dataset LLaVa1.5-7B Time
$\epsilon=\nicefrac{{2}}{{255}}$ $\epsilon=\nicefrac{{4}}{{255}}$
VEAttack COCO 10.8 7.1 51min38s
Flickr30k 10.7 6.3 49min59s
VEAttack+LLM COCO 12.3 7.3 3h11min7s
Flickr30k 10.1 5.4 3h8min58s
Appendix F More evaluation models and metrics on image captioning
Table 8 presents a comprehensive evaluation of the VEAttack method on image captioning across the COCO and Flickr30k datasets using the LLaVa1.5-7B model. Based on Eq. 5, we incorporate attacks from adversarial training algorithms (Mao et al., 2023) as proposed in the robust CLIP framework, as a baseline method to enable a more comprehensive comparison. The formulation can be expressed as $\mathcal{L}_{cls}:\ \tilde{v}=\mathop{\text{arg}\max}_{\|\delta\|\leq\epsilon}-\mathbb{E}_{j}\left[\cos\left(\tilde{z}_{cls},z_{text}^{j}\right)\right]$, where text features $z_{text}^{j}$ is extracted from user instructions. Compared to the clean baselines, VEAttack consistently demonstrates the most significant performance degradation on COCO captioning across all evaluation metrics. Specifically, $\text{BLEU}@4$, which focuses on measuring caption precision, shows a substantial drop of $86.7\%$ on the COCO dataset, indicating the highly effective nature of our VEAttack. ROUGE_L, which evaluates recall and sequence overlap, declines moderately $45.8\%$, suggesting that while the attack disrupts overall performance, some structural coherence in the captions remains. SPICE, focusing on semantic propositional content, drops $87.1\%$, indicating a broad degradation in meaningful content representation, reinforcing the effectiveness of VEAttack.
@@ Appendix F More evaluation models and metrics on image captioning
Datasets Attack $\text{BLUE}@4$ METEOR ROUGE_L CIDEr SPICE
@@
To demonstrate the broad applicability of VEAttack beyond CLIP-based architectures, we extend our evaluation to LVLMs utilizing diverse vision encoders. Specifically, we test MiniGPT-4 (Zhu et al., 2023), which employs the BLIP-2 (Li et al., 2023a) as a vision encoder, and mPLUG-Owl2 (Ye et al., 2024), which uses a distinct visual backbone, MplugOwl. We compare VEAttack against two gray-box methods, AttackVLM-ii (Zhao et al., 2023) and VT-Attack (Wang et al., 2024c). Following the setting of VT-Attack, we set the perturbation budget $\epsilon=8/255$, step size $\alpha=1/255$, and attack iterations $t=100$. As shown in Table 9, VEAttack consistently outperforms the baselines, achieving the most significant performance degradation on both models. On MiniGPT-4, VEAttack reduces the CIDEr score to 13.7, compared to 24.8 for VT-Attack. Similarly, on mPLUG-Owl2, our method achieves a CIDEr score of 65.1, significantly lower than the other gray-box methods. These results confirm that VEAttack effectively disrupts the visual semantic representation regardless of the underlying vision encoder architecture, validating its architecture-agnostic nature.
Table 9: Attack performance comparison on COCO Captioning against more LVLMs.
Models Attack $\text{BLUE}@4$ METEOR ROUGE_L CIDEr SPICE
MiniGPT-4 Clean 25.0 27.9 53.1 94.0 21.3
AttackVLM-ii Zhao et al. (2023) 5.8 13.5 30.0 5.4 5.4
VT-Attack Wang et al. (2024c) 16.3 21.9 44.5 54.9 14.2
VEAttack 5.1 12.3 30.5 10.0 4.5
mPLUG-Owl2 Clean 38.6 30.2 59.5 139.2 24.3
AttackVLM-ii Zhao et al. (2023) 20.6 21.0 46.8 71.6 13.8
VT-Attack Wang et al. (2024c) 17.7 19.2 43.9 59.8 12.5
VEAttack 12.5 15.8 38.8 40.4 8.3
Appendix G Attack Performance on Image-Text Retrieval task
G.1 Attack Performance under white-box of Vision Encoder
We evaluate VEAttack on image–text retrieval using CLIP‑ViT and compare against SGA (Lu et al., 2023) and DRA (Gao et al., 2024) on Flickr30k in Table 10. Following prior work, we use Attack Success Rate (ASR) as the primary measure of transferability. In this setting, we report R@1 and R@10 for both Text Retrieval (TR) and Image Retrieval (IR), where each R@k is interpreted as ASR@k, the fraction of adversarial queries whose ground‑truth counterpart does not appear among the top‑k retrieved results. As shown in the table, VEAttack attains competitive ASR, while being substantially more efficient, reducing time to 10min38s and memory to 3.13 GB. These results indicate that our vision‑encoder–only optimization delivers strong retrieval‑level attack efficacy with markedly lower computational cost.
Table 10: Comparison of SGA, DRA and VEAttack on the Image-Text Retrieval task.
Method TR R@1 TR R@10 IR R@1 IR R@10 Time Memory
SGA Lu et al. (2023) 100 99.9 100 99.98 14min17s 9.14GB
DRA Gao et al. (2024) 100 99.9 100 99.96 26min31s 10.31GB
VEAttack 99.88 99.19 99.81 98.54 10min38s 3.13GB
G.2 Transfer Attack Performance of Vision Encoders
We additionally evaluate the transfer attack performance following the transfer-attack methods, SGA (Lu et al., 2023) and DRA (Gao et al., 2024). We conduct transfer attacks using CLIP-ViT as the source model to attack ALBEF and CLIP-CNN. As shown in Table 11, while SGA and DRA are specifically designed with complex augmentation strategies to maximize transferability, VEAttack achieves competitive performance with fewer computational memory costs. Crucially, leveraging our Möbius Band in Observation 3, we replace the standard CLIP-ViT source with the robust FARE encoder. This simple substitution dramatically boosts VEAttack’s transferability, requiring no complex algorithm changes but achieving an R@1 of $35.35\%$ on ALBEF (Text Retrieval) and $48.78\%$ (Image Retrieval). This performance rivals the sophisticated DRA method while consuming only 10.52GB of memory. This demonstrates that utilizing robust vision encoders is a highly efficient and effective strategy for transfer attacks, offering a superior trade-off between performance and resource consumption.
Table 11: Comparison of transfer attack with SGA and DRA on the Image-Text Retrieval task.
Method Source ALBEF (TR R@1) ALBEF (IR R@1) CLIP-CNN (TR R@1) CLIP-CNN (IR R@1) Memory
SGA CLIP-ViT 22.42 34.59 53.26 61.1 16.63GB
DRA CLIP-ViT 27.84 42.84 64.88 69.50 16.71GB
VEAttack CLIP-ViT 16.58 32.23 44.06 53.55 6.75GB
FARE 35.35 48.78 55.17 63.05 10.52GB
Appendix H Attack Performance on Visual Grounding task
Table 12: Attack performance after difference methods on Visual Grounding task.
Method RefCOCO RefCOCO+
TestA TestB Val TestA TestB Val
Clean 49.0 53.9 48.2 50.8 49.1 47.3
AttackVLM-$i$ 27.6 33.5 28.9 28.5 27.9 29.1
VT-Attack 35.4 40.7 34.5 34.7 37.9 35.7
VEAttack 23.1 32.1 27.1 24.6 29.6 27.4
To verify the generalizability of VEAttack, we extended our evaluation to visual grounding task and select ReCLIP (Subramanian et al., 2022), a popular zero-shot baseline for visual grounding that relies on the CLIP vision encoder to localize objects. We evaluated the attack performance on the RefCOCO and RefCOCO+ datasets (Yu et al., 2016). For the attack settings, we maintain consistency across all comparison methods by setting the perturbation budget to $\epsilon=2/255$ and the number of iterations to $t=100$. As shown in Table 12, VEAttack exhibits effectiveness in disrupting the spatial reasoning capabilities. On the RefCOCO validation set, VEAttack degrades the performance to 27.1, significantly outperforming the baselines. This demonstrates that our vision-encoder-only attack effectively propagates to spatial localization tasks, further validating its downstream-agnostic nature.
Appendix I Transfer attack from larger specific vision encoders
We conducted transfer attacks using the vision encoders of mPLUG-Owl2 and Qwen-VL as source models on 1000 random images of COCO dataset with the CIDEr metric of the captioning task in Table 13 with $\epsilon=16/255,\alpha=16/(255\times 5)$. Surprisingly, we found that for both VEAttack and VT-Attack, when using the vision encoders of mPLUG-Owl2 and Qwen-VL as source models for transfer attacks, it is challenging to achieve successful attacks on other models. One potential reason could be that models tend to develop more specialized feature representations tailored to the specific LVLM architecture and training distribution as the scale of training data and network complexity increases. It might reduce the transferability of adversarial perturbations across diverse model architectures. In addition, the generated adversarial perturbations from these large vision encoders are concentrated in high-dimensional feature spaces. When transferred to smaller models such as CLIP, these perturbations may fail to effectively decouple or disrupt downstream task performance, as the limited capacity of smaller models may hinder their ability to process or adapt to such complex disturbances. This finding provides an empirical explanation for the community’s current practice of using an ensemble of foundational vision encoders like CLIP for attacks (Zhao et al., 2023; Zhang et al., 2024; Li et al., 2025; Jia et al., 2025).
Table 13: Transfer attack on COCO captioning from mPLUG-Owl2 and Qwen-VL to other models.
Method Source Target CIDEr Source Target CIDEr
VEAttack mPLUG-Owl2 mPLUG-Owl2 9.1 Qwen-VL Qwen-VL 19.3
Qwen-VL 131.8 mPLUG-Owl2 128.3
LLaVa1.5-7B 114.5 LLaVa1.5-7B 119.4
OpenFlamingo-9B 84.5 OpenFlamingo-9B 86.0
VT-Attack mPLUG-Owl2 mPLUG-Owl2 30.1 Qwen-VL Qwen-VL 55.3
Qwen-VL 132.1 mPLUG-Owl2 130.8
LLaVa1.5-7B 119.6 LLaVa1.5-7B 121.5
OpenFlamingo-9B 91.6 OpenFlamingo-9B 96.3
Figure 9: t-SNE visualization of visual feature distributions across different LVLMs.
Analysis on feature representation. To delve deeper into the underlying causes of the limited transferability observed in Table 13, we visualize the feature distributions of different vision encoders. Specifically, we randomly sample 100 patch tokens from the COCO validation set encoded by the vision encoders of LLaVa1.5-7B (CLIP-ViT-L), mPLUG-Owl2 (MplugOwl), and Qwen-VL (CLIP-ViT-bigG). As illustrated in Fig. 9, the t-SNE visualization reveals significant distributional discrepancies among these models. The feature spaces of these encoders form distinct, isolated clusters with large margins, indicating that the specific vision encoders used in large-scale LVLMs undergo substantial specific adaptation or fine-tuning, leading to a severe feature shift compared to standard foundation models like CLIP. Consequently, adversarial perturbations optimized on the specific manifold of one large encoder (e.g., Qwen-VL) are likely orthogonal or irrelevant to the feature space of other models (e.g., LLaVa).
Appendix J Evaluation on image classification datasets